OpenStudy (anonymous):

Solve for x. -5x+12=-12x+10

OpenStudy (anonymous):

IS THIS THE ANSWER TO MY QUESTION?

OpenStudy (anonymous):

x -> -(2/7)

OpenStudy (anonymous):

no;)

OpenStudy (amistre64):

what an odd response ....

OpenStudy (anonymous):

i was saying no to tarver..

OpenStudy (anonymous):

the answer is -2/7

OpenStudy (lilg132):

-5x +12 = -12x + 10 -5x + 12x = 10 - 12 7x = -2 7x / 7 = -2 / 7 remember to click good answer

OpenStudy (lilg132):

x = -2/7

OpenStudy (amistre64):

i was aware of that :) the odd part was Travers ....

OpenStudy (lilg132):

haha